40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 4 20:54:40 2020, gautam, Update, SUS, MC1 Slow Bias issues 
    Reply  Thu Jun 4 21:32:00 2020, Koji, Update, SUS, MC1 Slow Bias issues 
       Reply  Thu Jun 25 15:11:00 2020, gautam, Update, SUS, MC1 coil driver resistance quartered MCF.pdf
          Reply  Sun Jun 28 15:30:52 2020, gautam, Update, SUS, MC1 sat-box de-lidded 20200628T144138.jpg
             Reply  Sun Jun 28 16:29:58 2020, rana, Update, SUS, MC1 sat-box de-lidded 
                Reply  Sun Jun 28 17:36:35 2020, gautam, Update, SUS, MC1 sat-box de-lidded 
                   Reply  Mon Jun 29 11:55:46 2020, gautam, Update, SUS, MC1 sat-box de-lidded SatBoxSwap.jpg
                      Reply  Mon Jun 29 20:30:53 2020, Koji, Update, SUS, MC1 sat-box de-lidded 
Message ID: 15434     Entry time: Sun Jun 28 15:30:52 2020     In reply to: 15431     Reply to this: 15435
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 sat-box de-lidded 

Judging by the summary pages, some 18 hours after this change was made and the board re-installed, the MC1 shadow sensors began to report frequent glitches. I can't think of a plausible causal connection, especially given the 18 hour time lag, but also hard to believe there isn't one? As a result, the IMC is no longer able to stay locked for extended periods of time. I did the usual cable squishing, and also took off the lid to see if that helps the situation.

While the reduced series resistance means there is more current flowing through the slow path

  1. There isn't actually an increase in the net current flowing through the satellite box - this change just re-allocates the current from the fast path to the slow path, but by the time it reaches the satellite box, the current is flowing through the same conductor.
  2. afaik, the current buffers on the coil driver aren't overdriven - they are rated for 300 mA. No individual coil is drawing more than 30 mA.
  3. the resistors themselves should be running sufficiently below their rated power of 3W (I estimate 2.5 V ^2 / 100 ohms ~ 60 mW).
  4. The highest current should be through the UL and LR coils according to the voltage outputs from the Acromag. But the UL coil doesn't show significant glitching, and the LL one does despite drawing negligible DC current.

The attached FLIR camera image re-inforces what we already know, that the thermal environment inside the satellite box is horrible. The absolute temperature calibration may be off, but it was difficult to touch the components with a bare finger, so I'd say its definitely > 70 C.

Quote:

I implemented this change today. We only had 100 ohm, 3W resistors in stock (no 200 ohm with adequate power rating). Assuming 10 V is dropped across this resistor, the power dissipation is V^2/R ~ 1 W, so we should have sufficient margin. DCC entry has been updated with new schematic and photo of the component side of the board. Note that the series resistance of the fast actuation path was untouched.

Attachment 1: 20200628T144138.jpg  1.748 MB  Uploaded Sun Jun 28 16:38:38 2020  | Hide | Hide all
20200628T144138.jpg
ELOG V3.1.3-